# Atomic Smart Contract Interoperability with High Efficiency via Cross-Chain Integrated Execution

### v1

@@ VII-B3 Concurrency Performance with Fine-Grained State Lock
TABLE I: Gas consumption during cross-chain deployment.
@@
VII-B4 Logic-State Decoupling
We evaluate the gas consumption during the cross-chain smart contract deployment, using the Train-and-Hotel contracts as an example. We compare the results of the train and hotel contracts with and without logic-state decoupling. The experimental results in Table I indicate that, after applying logic-state decoupling, gas consumption during off-chain clone and deployment is reduced by 48.6% for the train contract and 74.5% for the hotel contract. The main reason is that, the train contract is implemented with more complex logic, while the hotel contract stores more states. As can been seen, for contracts with substantial state but relatively simple logic, logic-state decoupling can significantly reduce gas consumption.
@@ VIII Discussions and Limitations
Cross-Chain vs. Cross-Shard. Some existing works have explored the issue of cross-shard smart contract handling [58, 51]. However, cross-chain and cross-shard scenarios are fundamentally different, and their solutions cannot be directly applied to cross-chain contexts. The primary reasons are as follows: First, research on blockchain sharding typically involves modifications to the underlying system. In contrast, a key requirement of cross-chain protocols or systems is that they must not require modifications to the underlying blockchains, ensuring better compatibility with existing blockchain systems. Second, a blockchain sharding system usually has a beacon chain responsible for coordinating progress across shards, which is impractical in cross-chain protocols. After all, in cross-chain scenarios, each blockchain essentially belongs to a different system. Finally, in cross-chain contexts, blockchains are often heterogeneous (e.g., different consensus protocols), which is uncommon in blockchain sharding systems.

### v2

@@ VII-C3 Concurrency Performance with Fine-Grained State Lock
VII-C4 Logic-State Decoupling
We evaluate the gas consumption during the cross-chain smart contract deployment, using the Train-and-Hotel contracts as an example. We compare the results of the train and hotel contracts with and without logic-state decoupling. The experimental results in Table II indicate that, after applying logic-state decoupling, gas consumption during off-chain clone and deployment is reduced by 48.6% for the train contract and 74.5% for the hotel contract. The main reason is that, the train contract is implemented with more complex logic, while the hotel contract stores more states. As can been seen, for contracts with substantial state but relatively simple logic, logic-state decoupling can significantly reduce gas consumption.
TABLE II: Gas consumption during cross-chain deployment.
@@ VII-C4 Logic-State Decoupling
VII-C5 On-chain Operation During Off-chain Clone and Deployment
We evaluate the gas consumption and latency associated with the on-chain operations during the off-chain clone and deployment phase of the Hybrid Cross-Chain Smart Contract Deployment Protocol. The results are shown in Figure 10. Figure 10(a) presents the latency of different operations, which are nearly identical for both the Hotel and Train contracts. Because both the Deploy Logic and Register Contract operations complete within a single on-chain transaction, their latency remains low. Verification, by contrast, requires cross-chain validation and therefore incurs higher latency. Figure 10(b) shows the gas consumption of these operations. The Train contract consumes more gas for logic deployment due to its more complex functionality. However, for contract registration and verification, both contracts perform similar operations, leading to nearly identical gas usage. It is important to note that these operations are one-time costs. Overall, the results demonstrate that the on-chain overhead is reasonably lightweight and practical in terms of both gas and latency.
(a) Latency
(b) Gas consumption
Figure 10: Latency and gas consumption of different on-chain operations during off-chain clone and deployment.
@@ VIII Discussions and Limitations
Cross-Chain vs. Cross-Shard. Some existing works have explored the issue of cross-shard smart contract handling [59, 48]. However, cross-chain and cross-shard scenarios are fundamentally different, and their solutions cannot be directly applied to cross-chain contexts. The primary reasons are as follows: First, research on blockchain sharding typically involves modifications to the underlying system [60, 61]. In contrast, a key requirement of cross-chain protocols or systems is that they must not require modifications to the underlying blockchains, ensuring better compatibility with existing blockchain systems. Second, a blockchain sharding system usually has a beacon chain [62, 63] responsible for coordinating progress across shards, which is impractical in cross-chain protocols. After all, in cross-chain scenarios, each blockchain essentially belongs to a different system. Finally, in cross-chain contexts, blockchains are often heterogeneous (e.g., different consensus protocols), which is uncommon in blockchain sharding systems.